OpenStudy (anonymous):

why is any number to the power of 0 equal to 1?

OpenStudy (anonymous):

think of it like this: \[x^0 = x^{k-k} = x^k*x^{-k} = \frac{x^k}{x^{k}} = 1\]

OpenStudy (anonymous):

check out this link http://www.calculatoredge.com/math/mathlogic/logicans4.htm